OpenStudy (anonymous):

its 5x+12.5y=0 a direct variation if it is whats the constant

OpenStudy (anonymous):

Yes it is because: 5x + 12.5y = 0 12.5y = -5x y = (-5/12.5)x = -2/5x since every change in x results in a -2/5 change in y, the relationship is direct. the constant is -2/5 or -0.4

OpenStudy (anonymous):

thanks